# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

### v5

@@ Abstract
Method: Baseline recordings from a prior EEG study of 93 subjects, 37 with MS, were analyzed. Spectral bands served to compute cognetoms and categorize subsequent feature combination sets.
Result: A significant correlation arose between brain states predictors, clinical data and disease duration. Using cognetoms and spectral bands, a cross-sectional comparison separated patients from controls with a precision of 82% while using bands alone arrived at 64%.
Conclusion: Brain states analysis successfully distinguishes controls from patients with MS. The congruity with disease duration is a neurobiological indicator for disease accumulation over time. Our results imply that data-driven comparisons of EEG data may complement customary diagnostic methods in neurology and psychiatry. However, thinking ahead for quantitative monitoring of disease time course and treatment efficacy, we hope to have established the analytic principles applicable to longitudinal clinical studies.
@@ 1 Introduction
Mental illness is an enormous burden for societies worldwide (Smith, 2011; Vigo et al., 2016). Above all it is the decay of cognitive functioning, the essence of human reasoning, that turns a mental illness into a devastating and tangible experience for patient and surrounding alike. Nonetheless diagnosis and staging of diseases still rely on customary clinical, laboratory and imaging tests by which cognitive integrity is not assessed on biological grounds. Electrical scalp recordings exist for a century and were used decades ago as diagnostic evidence in multiple sclerosis although the analytic methods lacked the quantitative aspect for relating brain measures to disease time course or treatment outcome. We believe that a modern data-driven functional brain mapping analysis approach using non-invasive EEG data will meaningfully enhance the prognostic power for clinical trials and facilitate the routine assessment of drug efficacy and safety through objective measures for a changing cognito-electrical substrate.
@@
The central nervous system (CNS) is an electrical organ and invokes with each mental act a plethora of neural activities instantly reflected in the electrical brain space and recordable on the skull surface with electrode arrays (Buzsáki, 2005). Electrical states of brain functioning, coined cognetoms, change at any given moment (Koenig et al., 1999; Samdin et al., 2017) and are traced with physiological measures while anatomical brain lesions evolve slower over days or weeks and are monitored with MRI exams (Eshaghi et al., 2021; Tommasin et al., 2021; Lazzarotto et al., 2024). The crucial difference is in the immediacy and temporal accuracy of electrical measures and the dispersed soft demarcation of electrical phenomena on the skull surface that mirror underlying distributed brain activity. These diffuse and perpetually evolving cloud-like arrangements are due to oscillating neural networks and reflect equally housekeeping and cognitive neural processes (Buzsáki, 2005).
@@
MS is a chronic, relapsing-remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
@@
Recent advances in pharmaceutical research for aggressive disease-modifying treatments lead to a surge of novel drugs for MS over the last two decades. Currently, there are 18 disease-modifying therapies approved by Health Canada (relapsing remitting form) while heaps of new compounds are being tested in clinical trials for clinical and MRI outcomes. The desired effect of drug therapy on cognitive preservation (Falet et al., 2022) however is still conventionally assessed with questionnaires and neuro-psychological tests although a recent upswing of publications using functional brain measures in MS is observable (Kisler et al., 2020; Paolicelli et al., 2021; Santinelli et al., 2021; Mohseni and Moghaddasi, 2022; Sorrentino et al., 2022).
We adopted the statistical principles of our former brain states analysis method (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013) and tested the new method on data taken from a prior EEG study about multiple sclerosis in healthy subjects and patients (Jouzizadeh et al., 2021). Our motivation to analyze such cross-sectional cohort data roots in the current lack of access to serial EEG measures in MS patients. While our methods are primarily developed for the purpose of comparing serial EEG measures longitudinally within the same individual we felt confident that the data modeling phase could also generate robust enough brain state maps from single-time-point data to meaningfully support a cross-sectional comparison of brain states features between two subject cohorts. We show here the tangential application of a brain states EEG analysis method that successfully distinguishes healthy data from MS data.
@@
Figure 1: (A) Result of dimension reduction with un-supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual, red color with MS patients, blue with control subjects. (B) Result of dimension reduction with supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual. We split the data into train (60%) and test (40%) subsets of the data. Green corresponds with the train subset of control subjects and blue with the test subset, then orange with train subset of MS patients and red with the test subset. In both plots, there is no meaning to axis orientation and values, only the relative position of points in two-dimensional space is relevant as well as the density distribution through clustering.
@@ 2.1 Subjects and EEG recording
The EEG data of 93 subjects were taken from an earlier project where participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 37 subjects diagnosed with MS (24 female, 13 male, mean age 29.5 years, age range 20-53 years) and 56 healthy volunteers (29 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project. The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
@@ 2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files sans further metadata or personal information. Electrode labels and anatomical positional information were discarded and played no further role during data analysis. Attention was paid to maximally preserve the fine-grained intrinsic physiological information present in the data (Meisler et al., 2019) for the brain states generation phase further downstream in the processing pipeline. For these reasons EEG data were taken as is without applying common data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and without introducing pre-processing related artifacts (McDermott et al., 2021). This avenue is possible because the data processing methods during brain state detection and later analysis phases intrinsically suppress noise components in the data throughout the entire brain states processing pipeline (Janoos et al., 2011; Mórocz et al., 2012) and render the approach fundamentally apt and robust in dealing with clinical-grade real-life data (Janoos et al., 2013). This ‘automatic noise robustness’ lies in the fact that brain states detected throughout the data sets bear inter-individual characteristics and are hence not prone to include meaningless individual noise properties not universally present across the entire data set.
@@ 2.3 Brain states detection (BSD)
We define brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals. Brain states are understood as moments of distributed brain activity of a given time point (Koenig et al., 1999; Janoos et al., 2011; Mórocz et al., 2012). In a first move a limited set of eight brain states are computed. These sets serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level. EEG recordings were segmented into data epochs where each individual epoch, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz) that are extracted from raw EEG data.
@@
In a second move the best match distribution of brain states is detected for each time point along the EEG data time axis with weights shown for each brain state. In other words, our approach translates the multi-channel raw EEG signal of each subject into a sequence of brain states as qualitative feature with the associated weights as numerical quantitative feature.
@@ 3 Results
3.1 Brain states analysis (BSA)
In order to capture differences between the EEG data of healthy persons and patients with MS two approaches were applied : manifold learning and catboost classifier. For this, 80 distinct feature combinations were built and are described as combinations of 8 brain state weights, two experimental recording conditions (eyes-open and eyes-closed) and five EEG frequency band amplitudes as mentioned.
The first approach, manifold learning, reduces the dimensionality of the individual data (McInnes et al., 2018). The previously prepared sequence of brain states and weights are used to embed the individual data into a two-dimensional space. Each point in that space corresponds with one person. First we applied the non-supervised version of the UMAP algorithm for which the results reveal a distinct trend separating control from MS cases (upper figure 1). Next we applied the supervised version by splitting data into train (60%) and test (40%) subsets of data. The trained model produced the two-dimensional embedding with high separability between the two subject groups (lower figure 1).
Our second approach used the catboost classifier method (Dorogush et al., 2018) which first splits the data into train (60%) and test (40%) subsets of data, then trains the model, and last evaluate the quality of the model on the test data subset. The quality of the model is described by precision and F-score values.
In sum, our very first implementation of the brain states approach for EEG analysis shows a successful performance for detecting MS in common clinical low-resolution EEG data sets at a high precision of 82%.
separation based on precision F-score
frequency bands alone 0.64 0.61
frequency bands and brain states 0.82 0.78
Table 1: Performance of MS detection with catboost classifier. The inclusion of brain states features augments significantly the precision for distinguishing EEG data sets in patients with MS vs healthy subjects as compared to taking into account the frequency band features alone.
@@ 3.2 Benefit of the brain states approach
We demonstrate a direct benefit for including brain states features in the analysis as compared to taking into account only the frequency band features typically used for EEG analysis (table 1). A benchmark for the two EEG analysis approaches show an impressive performance boost of 18% for distinguishing EEG data sets of patients with MS versus healthy subjects in favor of the combined method. The precision for detecting MS using frequency bands alone reaches 64% while adding information about brain states arrives at 82%.
Figure 2: Specific feature analysis for MS detection. The relative importance of feature combination (brain states vs frequency bands vs experimental conditions) is shown. (A) The upper figure part depicts ordered feature combinations (x-axis) vs feature importance (y-axis). (B) The lower part shows the comparison between the control vs the patient group regarding feature importance. The blue line stands for the control group, the orange line for MS. Only the 40 top most important feature combinations are shown.
@@ 3.3 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two subject groups into healthy vs MS is outlined in figure 2. We observe that a set of certain feature combinations that have a high relative importance, include both eyes-open and eyes-closed conditions and the delta frequency band appear to drive the separation between healthy and MS subjects.

### v6

@@ Abstract
Method: Baseline recordings from a prior EEG study of 88 subjects, 36 with MS, were analyzed. Spectral bands served to compute cognetoms and categorize subsequent feature combination sets.
Result: The brain states predictor correlates with disease burden and duration. Using cognetoms and spectral bands, a cross-sectional comparison separated patients from controls with a precision of 85% while using bands alone arrived at 79%.
Conclusion: We demonstrate the efficiency of the quantitative data-driven method based on brain states analysis by contrasting EEG data of patients with MS and healthy subjects. The congruity with disease severity and duration is a neurophysiological indicator for disease accumulation over time. We discuss potential applications of the approach for the monitoring of disease time course and treatment efficacy in longitudinal clinical studies in psychiatry and neurology.
@@ 1 Introduction
Mental illness is an enormous burden for societies worldwide (Smith, 2011; Vigo et al., 2016). Above all it is the decay of cognitive functioning, the essence of human reasoning, that turns a mental illness into a devastating and tangible experience for patient and surrounding alike. Nonetheless diagnosis and staging of diseases still rely on customary clinical, laboratory and imaging tests by which cognitive integrity is not assessed on biological grounds. Electrical scalp recordings have existed for a century and have been used decades ago as diagnostic evidence in multiple sclerosis, although the analytic methods lacked the quantitative aspect needed to relate brain measures to disease time course or treatment outcome. We believe that a modern data-driven functional brain mapping analysis approach using non-invasive EEG data will meaningfully enhance the prognostic power for clinical trials and facilitate the routine assessment of drug efficacy and safety through objective measures for a changing cognito-electrical substrate.
@@
The central nervous system (CNS) is an electrical organ and invokes with each mental act a plethora of neural activities instantly reflected in the electrical brain space and recordable on the skull surface with electrode arrays (Buzsáki, 2005). Electrical states of brain functioning, coined cognetoms, change at any given moment (Koenig et al., 1999; Samdin et al., 2017) and are traced with physiological measures while anatomical brain lesions evolve slower over days or weeks and are monitored with MRI exams (Eshaghi et al., 2021; Tommasin et al., 2021; Lazzarotto et al., 2024; Goebl et al., 2025). The crucial difference is in the immediacy and temporal accuracy of electrical measures and the dispersed soft demarcation of electrical phenomena on the skull surface that mirror underlying distributed brain activity. These diffuse and perpetually evolving cloud-like arrangements are due to oscillating neural networks and reflect equally housekeeping and cognitive neural processes (Buzsáki, 2005).
@@
MS is a chronic, often relapsing-remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025; Pauletti et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS, with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
@@
Recent advances in pharmaceutical research for aggressive disease-modifying treatments lead to a surge of novel drugs for MS over the last two decades. Currently, there are 18 disease-modifying therapies approved by Health Canada (relapsing remitting form), while heaps of new compounds are being tested in clinical trials for clinical and MRI outcomes. The desired effect of drug therapy on cognitive preservation (Falet et al., 2022), however, is still conventionally assessed with questionnaires and neuro-psychological tests, although a recent upswing of publications using functional brain measures in MS is observable (Kisler et al., 2020; Paolicelli et al., 2021; Santinelli et al., 2021; Sarrias-Arrabal et al., 2021; Mohseni and Moghaddasi, 2022; Sorrentino et al., 2022).
For this work we propose an extension for the concept of brain states analysis (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013). Instead of the statistical principles based on hidden Markov model (HMM) and functional MRI data we developed a novel method for brain states detection and analysis of EEG signals taken from a prior cross-sectional EEG study about patients with multiple sclerosis and healthy subjects (Jouzizadeh et al., 2021).
@@
Figure 1: (A) Result of dimension reduction with un-supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual, red color with MS patients, blue with control subjects. (B) Result of dimension reduction with supervised version of UMAP manifold learning algorithm. The sequence of brain states for each individual data is embedded to a two-dimensional space. Each point corresponds to one individual. Green corresponds with the train subset of control subjects and blue with the test subset, then yellow with train subset of MS patients and red with the test subset. In both plots, there is no meaning to axis orientation and values, only the relative position of points in two-dimensional space is relevant as well as the density distribution through clustering.
@@ 2.1 Subjects and EEG recording
The EEG data of 88 subjects were taken from an earlier project. Participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 36 subjects diagnosed with MS (23 female, 13 male, mean age (29.6 years, age range 20-53 years) and 52 healthy volunteers (25 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project (see tables in Appendices A and B for subject information). The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
@@ 2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files without further metadata or personal information. Electrode labels and anatomical positional information were discarded and played no further role during data analysis. Attention was paid to maximally preserve the intrinsic physiological information present in the data for the brain states generation phase further downstream in the processing pipeline. EEG data were taken as is without applying common manual or automated data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and hence without introducing data pre-processing related artifacts.
The need for artifact removal is an ongoing debate in the field. For example, in stroke patients who undergo real-time EEG-based brain computer interface (BCI) therapies, the paradigm-related muscular activity in naturalistic movements can carry a significant informative and therapeutic value even as artifactual non-brain-derived components. In contrast, the very same components may detrimentally affect analysis of brain physiological data (McDermott et al., 2021). Another example where noise filtering can remove a mixture of both artifactual and valid signal is found in intracranial EEG studies because artifacts linked just loosely to behavior remain poorly characterized (Meisler et al., 2019). Instead the authors suggest to increase the number of observations and so to enhance statistical power.
Resting state paradigms come with a random pattern of incidental and intentional muscular activities that surface as subject specific chance noise in space and time. The data processing pipeline during the brain state detection and the later analysis phases intrinsically suppress noise components (Janoos et al., 2011; Mórocz et al., 2012) and fundamentally render the approach apt to handle clinical-grade real-life data (Janoos et al., 2013). This ‘design-specific noise robustness’ lies in the fact that brain states are generated for the entire data set and hence hold inter-individual, universally present features but suppress intra-individual periodic or permanent signal properties. This way EEG artifacts due to motor, sensory and other effects do nor linger in brain states nor lead to subgroups in the analysis. Likewise is the subsequently computed disease predictor space not sensitive to patient-specific muscle artifacts in the electrical signal.
@@ 2.3 Brain states detection (BSD)
We were inspired by the similarity of a brain states analysis method we had developed for processing rapid functional MRI data (Janoos et al., 2011; Mórocz et al., 2012) and the concept of microstates analysis for EEG data introduced by Lehmann et al (Lehmann et al., 1987). Both approaches aim to transform the complex original data into a sequence of discrete, quasi-stable meaningful patterns called brain states or microstates. It was shown that the distribution of these states play a role as biomarkers in various psychiatric and neurological conditions including MS (Koenig et al., 1999; Gschwind et al., 2016).
We define brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals also understood as moments of distributed brain activity of a given time point (Janoos et al., 2011; Mórocz et al., 2012). In other words, states are defined as patterns of statistical dependencies or spatio-temporal correlations between brain regions measured across EEG electrodes. In this paper, a brain state is a typical configuration of spatial activity without actual knowledge of spatial coordinates of electrodes. The states serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level.
In a first move a limited set of eight brain states are computed. Raw EEG signals from each channel are transformed into sequences of overlapping spectrograms using the Fast Fourier Transform (FFT) implementation from the cusignal.spectrogram library (Thompson and Nicely, 2021). The recordings are segmented into data epochs where each individual window, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz). While window size and number of states are varied as technical optimization against detection performance, its length of 2s approximates the scan acquisition times needed for a single multi-slice T2* weighted brain shot typical for functional MRI studies (Mórocz et al., 2012; Janoos et al., 2013). To reduce dimensionality and capture coherent spatial patterns, we grouped EEG channels based on the similarity of their time series. Pairwise distances between channels were computed, followed by hierarchical clustering to identify groups of similar channels. For the feature vector construction, we computed at each time step the mean signal intensity across predefined frequency bands within each cluster of grouped channels. This resulted in a concatenated vector representation summarizing the spatio-spectral profile of the EEG activity at each moment. For the identification of brain states, we applied k-means clustering to the resulting vectors to identify a set of prototypical spatio-spectral configurations, which we interpret as data-driven representations of transient brain states.
@@
In a second move, for the temporal coding of brain states, the best match distribution of brain states is detected for each time point along the EEG data time axis with weights shown for each brain state. Each time step in the original sequence is assigned to its nearest brain state cluster, producing a simplified temporal sequence of discrete brain state codes, each associated with signal intensities across frequency bands. In other words, our approach translates the multi-channel raw EEG signal of each subject into a sequence of brain states as qualitative feature with the associated weights as numerical quantitative feature.
2.4 Brain states analysis (BSA)
In order to capture differences between the EEG data of healthy persons and patients with MS two approaches were applied : manifold learning and catboost classifier. For the subject-level feature extraction, 80 distinct feature combinations were built and are described as combinations of 8 brain state weights, two experimental recording conditions (eyes-open and eyes-closed) and five EEG frequency band amplitudes as mentioned. This compressed representation enabled further statistical analysis. For each subject, we computed distributions over brain states and average signal intensities per state. We used a fixed-length feature vector for each subject, defined by the combination of brain states (8), frequency bands (5), and experimental conditions (2), for the downstream comparative analyses.
The first approach, manifold learning, reduces the dimensionality of the individual data (McInnes et al., 2018). The previously prepared sequence of brain states and weights are used to embed the individual data into a two-dimensional space. Each point in that space corresponds with one person. First we applied the non-supervised version of the UMAP algorithm for which the results reveal a distinct trend separating control from MS cases (upper figure 1). Next we applied the supervised version by splitting the data into train (60%) and test (40%) subsets. The trained model produced the two-dimensional embedding with high separability between the two subject cohorts (lower figure 1).
Our second approach used the catboost classifier model (Dorogush et al., 2018) which applied a 30-fold cross-validation to split the data into train (60%) and test (40%) subsets, then trains the model, and last evaluate the quality of the model on the test data subset. Only the binary labels for MS and non-MS were used for model training. All other clinical or demographic information was taken into account only for the post-analysis procedures. The quality of the model is described by precision, recall, f1-score and support (see table 1).
@@ 3.1 Benefit of the brain states approach
Our very first implementation of the brain states approach for EEG analysis shows a successful performance for detecting MS in common clinical low-resolution EEG data sets at a high precision of 85%. We demonstrate a direct benefit for including brain states features in the analysis as compared to taking into account only the frequency band features typically used for EEG analysis (table 1). A benchmark for the two EEG analysis approaches show a performance boost of 6% for distinguishing EEG data sets of patients with MS versus healthy subjects in favor of the combined method. The precision for detecting MS using frequency bands alone reaches 79% while adding information about brain states arrives at 85%.
separation based on precision recall f1-score support
frequency bands alone 0.79 0.78 0.78 36
frequency bands & brain states 0.85 0.84 0.84 36
Table 1: Performance of MS detection with catboost classifier. We compare the classifier performance for two scenarios : first the features are defined as frequency bands and experimental conditions such as eyes open and eyes closed. In the second scenario the features are defined as a combination of brain states, frequency bands and experimental conditions. We applied a 30-fold cross-validation to split the data into train (60%) and test (40%) subsets of the data. Taking into account the brain states augments the precision, recall and f1-score for distinguishing EEG data sets in patients with MS vs healthy subjects - as compared to taking into account the frequency band features alone.
@@ 3.2 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two cohorts into healthy vs MS is outlined in figure 2. We observe that a set of feature combinations of high importance include both eyes-open and eyes-closed conditions and the delta frequency band that appear to drive the separation between healthy and MS subjects. Averaged frequency spectra for both cohorts are shown as graphs in Appendix C.
Figure 2: Specific feature analysis for MS detection. The relative importance of feature combination (brain states vs frequency bands vs experimental conditions) is shown. For example the first feature combination is called ‘Delta_5_ec’ which stands for ‘Delta’ frequency band, brain states no. ‘5’, and ‘ec’ for the eyes closed condition. (A) The upper figure part depicts ordered feature combinations (x-axis) vs feature importance (y-axis). (B) The lower part shows the comparison between the control vs the patient cohort regarding feature importance. The blue line stands for the control group, the orange line for MS. Only the 40 top most important feature combinations are shown.